OpenStudy (anonymous):

Find the approximate value of the circumference of a circle with the given radius. Use = 3.14. Round your results to one more decimal than in the given radius. 4 inches

geerky42 (geerky42):

\[C = 2\pi r\] Just plug in value and solve for C.

OpenStudy (anonymous):

how do i do that where do i plug it in

geerky42 (geerky42):

You plug 4 in r because r (radius) is 4 inches

geerky42 (geerky42):

for \(\pi\), you just use 3.14 instead

OpenStudy (anonymous):

nd what do i do with the 3.14

geerky42 (geerky42):

multiply it with 4 and 2

geerky42 (geerky42):

C = 2 * 3.14 * 4

OpenStudy (anonymous):

so i multiply 2 times 3.14 times 4

geerky42 (geerky42):

right

geerky42 (geerky42):

multiply 2 times 3.14 times 4

OpenStudy (anonymous):

25.12?

geerky42 (geerky42):

none. that's 2 times 4. you need to do 3.14 too

geerky42 (geerky42):

what is 2 times 3.14?

geerky42 (geerky42):

Sorry, my internet is horrible. Yes it's 25.12 25.12 is your answer.

OpenStudy (anonymous):

thats not an answer choice :(

OpenStudy (anonymous):

12.5 inches 12.6 inches 25.1 inches 25.2 inch